OpenStudy (anonymous):

What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)?

OpenStudy (anonymous):

see u can write y=x+7 slope of this line is 1 for perpendicularity condition is: (slope of line 1)*(slope of line 2)=-1 hence slope of other line comes out to be -1 now apply (y-y1)=slope*(x-x1) (y+1)=-1*(x+1) so eqn of line is x+y+2=0 cheers :)